The circumference of the wheel of the tricycle of Timothy is 1/4 of the wheel of the car from his father. If the circumference of the wheel of his father is 16 π (pi), which is the radius of the wheel of Timothy?
Answer by edjones(8007)

t=timmy
C(t)=16pi/4=4pi=pi*d
r(t)=d/2=4/2=2
.
Ed